The average of cholestorel content a certain canned goods is 215 miligrams,and the standarddeviation is 15 miligrams.Assume that the variable is normally distributed.If a sample of 25 canned goods is selected,what is the probability that the mean of the sample will be than 220 miligrams?

1
Expert's answer
2022-04-07T07:08:36-0400

mean =215

standard deviation=15

n=25

x=220

P(X>220)=1-P(x"\le"220)

Z="\frac{x-mean}{\frac{s}{\sqrt n}}"

Z="\frac{220-215}{15/5}"

Z="\frac{5}{3}"

P(Z"\le 1.67" )=0.9525

P(X>220)=1-0.9525

P(X>220)=0.0475

The probability that the mean of the sample will be than 220 milligrams is 0.0475.